Jack collects the new nickels and quarters. The face value of he collection is $26.25. He has 50 more quarters than twice the number of nickels. How many nickels and how many quarters does he have?
Answer by jrb1965(16)

So you set up equations to match the words:
N = Nickels
Q = Quarters
50 more means "add 50"
Twice means "multiply by 2"
So Q = 2N + 50
For the value, $26.25 = $0.25Q + $0.05N, since quarters are $0.25 and nickels are $0.05
Substitute the equation for Q into the value equation and solve
$26.25 = $0.25(2N + 50) + $0.05N (multiply $0.25 by 2N and 50)
$26.25 = $0.50N + $12.50 + $0.05N (subtract $12.50 from both sides)
$13.75 = $0.50N + $0.05N (combine N terms)
$13.75 = $0.55N (divide by $0.55)
25 = N
Q = 2N + 50 = 2(25) + 50 = 50 + 50 = 100
Answer -
N = 25
Q = 100